-9y-5-(y-3). Simplify and show all your work please. thank you
Answer by Aztec(6)

-9y-5-(y-3)
Let's solve this bit by bit. -(y-3)=-y - - 3
= -y + 3<==== two minus signs would equal a plus sign.
So we substitute "-(y-3)" with "-y + 3"

-9y-5-y+3= -10y - 2